# Saliency-Aware Quantized Imitation Learning for Efficient Robotic Control

### v1

@@ 1 Introduction
Figure 1: Comparison of action discrepancy (L2-norm) between the quantized and full-precision (FP) policies in OpenVLA using the PTQ, QAT, and SQIL, and evaluation of the average success rate on the LIBERO benchmark.
@@ 4.1 Saliency-based State-Importance Score
Figure 4: Comparison of SIS and keyframe (KF) on trajectory.
4.2 Quantization-Robust Action Distillation
where $\alpha_{t}=\beta$ if $SIS^{s_{t}}_{\pi^{\text{FP}}}>T$, and $\alpha_{t}=1$ otherwise; $D$ is a discrepancy metric (e.g., L2-norm), and ${\pi}(s_{t})$ is the probability distribution of all possible actions for state $s_{t}$. The hyperparameter $\beta(>1)$ applies extra weight to high-importance states, and $T$ is a threshold that selects the top $p$=20of SIS values. Unlike conventional KD methods, QRD employs the weighting $\alpha_{t}$ to amplify the loss for states identified by SIS (i.e., $SIS^{s_{t}}_{\pi^{\text{FP}}}>T$). Training convergence is not sensitive to the hyperparameters $D(),\beta,T$, and we use the same values (details in Sec. 5.2.3) for all the experiments across robot control, self-driving, and physics simulation.
LIBERO Benchmark QAT SQIL (KF) SQIL (SIS)
Avg. Success Rate (%) 71.6 ± 0.5 72.6 ± 0.7 73.2 ± 0.6
Table 1: Saliency metric comparison: Keyframe (KF) and SIS.
4.3 Saliency-Aware Quantized Imitation Learning
@@ 4.4 Analysis
  Task #Trials FP #Successes SQIL W4 #Successes
Pick up the black bowl on the wooden cabinet and place it on the plate 50 34 29
Grab the black bowl from the wooden cabinet and set it on the plate 50 25 28
Remove the black bowl from the wooden cabinet and place it on the plate 50 31 34
@@
  Light Success Rate % $\uparrow$
Intensity FP SQIL W4
  100% 50.7% 49.2%
80% 50.5% 49.1%
60% 49.5% 48.8%
@@
  Method $D$ Suc. Rate $\uparrow$
SQIL L2-norm 73.2%
KL divergence 72.9%
  $\beta$ 1.5 2 3 4 5
Suc. Rate $\uparrow$ 49.0% 49.2% 48.9% 48.2% 47.8%
  $p$ 10% 15% 20% 30%
Suc. Rate $\uparrow$ 48.8% 49.3% 49.2% 48.5%
@@ 5.1 Experimental Settings
Robot Manipulation: We utilized OpenVLA [27] with weight quantization through AWQ [34] and QuaRot [13] methods. Evaluation were made using the the LIBERO [35] benchmark in simulated environments, spanning tasks like LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long. Each task was repeated in three rounds of evaluations. We fine-tuned a minimal set of trainable parameters (110M parameters) using SQIL and QLoRA (r=32) and tested it in a real-world scenario, replicating BridgeDataV2 [51], generating a dataset $\mathcal{D}_{\text{E}}$ of 80 episodes. This dataset was then used to refine OpenVLA with SQIL and QLoRA for deployment on the UR5 robot.
@@ 5.2.2 Generalization
Fig. 7 compares the performance results of various quantization techniques in real-world scenarios. Relative to the FP policy, the performance of PTQ significantly declines, while QAT exhibits slight improvements. Notably, SQIL achieves performance comparable to the FP policy, demonstrating that reducing quantization errors in important states directly contributes to achieving goals in real-world environments.
5.2.3 Ablation Study
Hyperparameter Sensitivity Analysis. We examine the sensitivity of hyperparameters within the SQIL, as detailed in Table 5. Opting for the L2-norm as the discrepancy metric $D$ yields slightly better results than the KL divergence. For SQIL, the hyperparameter $\beta$ shows stable performance within the 1.5 to 3 range. The threshold $T$ demonstrates resilience, delivering superior performance from the top 10% to 20% range, and maintaining robust results even at 30%, compared to QAT. These results confirm the robustness of our framework’s hyperparameters.
@@ 1.2 Hyperparameter Setting
Table 10: Comparison of success rates across different tasks in LIBERO-Spatial benchmark, evaluated on OpenVLA with INT4 quantization.
@@ 3 Ablation Study
4 Technical Clarifications
Action Discrepancy (Fig. 1) and Mission-Critical States. Fig. 1 highlights our key finding: INT4 policies fail abruptly during physical interaction with objects (e.g., picking, pulling, or closing). To analyze, we compute joint-normalized L2 action errors between FP and INT4 policies. Each episode is segmented into four sub-tasks, with timelines realigned only after a successful segment. We refer to the quantized policies’ failure-prone interaction points as mission-critical states, in contrast to routine ones along the trajectory.
Cost for SIS and Finetuning. SIS is a one-time process per existing expert dataset. We exploit spatio-temporal redundancy to reduce offline cost: perturbations are applied over a grid of $N\!\times\!N$ patches that evenly divide the image, and SIS is computed every $k$-th frame, reusing previous scores for intermediate frames. Empirically, $N\!=\!8$ and $k\!=\!4$ yielded robust results across all experiments and domains. Additionally, SQIL serves as a drop-in replacement for unavoidable fine-tuning in VLA domain adaptation [2] due to variations (e.g., camera viewpoints, lighting, and hardware) using the same expert data and hyperparameters.
Generalization: Model Distillation and Real-World Tasks: To demonstrate SQIL’s broad applicability, we applied it to $\pi_{0}$ [2], a compact and high-performant policy via model distillation. As shown in Table 12, SQIL consistently outperforms PTQ and QAT, achieving results near the FP baseline while delivering 3.0×3.0$\times$3.0 × memory savings. In response to the request for broader real-world scene coverage, we extended the setup from the ToySink enclosure to a tabletop workspace allowing larger motions. We collected 80 episodes and fine-tuned OpenVLA for three manipulation tasks. Table 13 shows that SQIL maintained FP-level performance.
LIBERO Benchmark FP PTQ QAT SQIL
Model Size 3.3B 1.1B 1.1B 1.1B
Avg. Success Rate 93.8% 92.0% 92.6% 93.3%
Table 12: INT4 quantization results for $\pi_{0}$ (model distillation).
  Task #Trials #Successes
FP PTQ QAT SQIL
(a) 30 25 22 25 24
(b) 30 24 20 22 24
(c) 30 22 18 17 21
               Suc.Rate(%) 79 67 71 77
Table 13: Real-world evaluation with OpenVLA on UR5.
Hyperparameter Robustness. As shown in Table 14, SQIL exhibits remarkable stability across hyperparameter variations, enabling shared settings across domains. Changing saliency noise distribution had negligible effect, while using raw SIS values as weights degraded performance—confirming the importance of our balanced weighting approach.

### v2

@@ 1 Introduction
Figure 1: Comparison of action discrepancy (L2-norm) between the quantized and full-precision (FP) policies in OpenVLA using the PTQ, QAT, and SQIL, and evaluation of the average success rate on the LIBERO benchmark. For fair comparison, episode timelines are segmented into four sub-tasks and realigned based on successful transitions.
@@ 4 Method
Figure 4: Comparison of SIS and keyframe (KF) on trajectory.
LIBERO Benchmark QAT SQIL (KF) SQIL (SIS)
Avg. Success Rate (%) 71.6 ± 0.5 72.6 ± 0.7 73.2 ± 0.6
Table 1: Saliency metric comparison: Keyframe (KF) and SIS.
@@ 4.1 Saliency-based State-Importance Score
4.2 Quantization-Robust Action Distillation
where $\alpha_{t}=\beta$ if $SIS^{s_{t}}_{\pi^{\text{FP}}}>T$, and $\alpha_{t}=1$ otherwise; $D$ is a discrepancy metric (e.g., L2-norm), and ${\pi}(s_{t})$ is the probability distribution of all possible actions for state $s_{t}$. The hyperparameter $\beta(>1)$ applies extra weight to high-importance states, and $T$ is a threshold that selects the top $p$=20of SIS values. Unlike conventional KD methods, QRD employs the weighting $\alpha_{t}$ to amplify the loss for states identified by SIS (i.e., $SIS^{s_{t}}_{\pi^{\text{FP}}}>T$). Training convergence is not sensitive to the hyperparameters $D(),\beta,T$, and we use the same values (details in Supplementary Sec. 3) for all the experiments across robot control, self-driving, and physics simulation.
@@ 4.4 Analysis
                                                                Language Instruction Generalization         Illumination Generalization
  Task #Trials # Successes $\uparrow$ Light Success Rate(%) $\uparrow$
FP SQIL W4 Intensity FP SQIL W4
  Pick up the black bowl on the wooden cabinet and place it on the plate 50 34 29 100% 50.7 49.2
Grab the black bowl from the wooden cabinet and set it on the plate 50 25 28 80% 50.5 49.1
Remove the black bowl from the wooden cabinet and place it on the plate 50 31 34 60% 49.5 48.8
@@ 5.1 Experimental Settings
Robot Manipulation: We utilized OpenVLA [27] with weight quantization through AWQ [34] and QuaRot [13] methods. Evaluation were made using the the LIBERO [35] benchmark in simulated environments, spanning tasks like LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long. Each task was repeated in three rounds of evaluations. We fine-tuned a minimal set of trainable parameters (110M parameters) using SQIL and QLoRA (r=32) and tested the model in two real-world scenarios: a replication of BridgeDataV2 [51] and a tabletop setup that enables broader manipulation behaviors. A dataset $\mathcal{D}_{\text{E}}$ of 80 episodes was collected and used to further refine OpenVLA with SQIL and QLoRA for deployment on the UR5 robot.
@@ 5.2.1 Full-Benchmark Evaluation
  Task #Trials #Successes
FP PTQ QAT SQIL
(a) 30 25 22 25 24
(b) 30 24 20 22 24
(c) 30 22 18 17 21
               Suc.Rate(%) 79 67 71 77
Figure 8: Real-world evaluation with OpenVLA on UR5.
LIBERO Benchmark FP PTQ QAT SQIL
Model Size 3.3B 1.1B 1.1B 1.1B
Avg. Success Rate 93.8% 92.0% 92.6% 93.3%
Table 5: INT4 quantization results for $\pi_{0}$ with OpenVLA.
@@ 5.2.2 Generalization
To further assess the generality of SQIL, we applied it to $\pi_{0}$ [2], a compact and high-performance policy obtained via flow matching-based action generation. As shown in Table 5, SQIL consistently outperforms PTQ and QAT, achieving performance close to the FP baseline while providing 3.0$\times$ memory savings.
Fig. 7 and Fig. 8 compares the performance results of various quantization techniques in real-world scenarios. Relative to the FP policy, the performance of PTQ significantly declines, while QAT exhibits slight improvements. Notably, SQIL achieves performance comparable to the FP policy, demonstrating that reducing quantization errors in important states directly contributes to achieving goals in real-world environments.
@@ 1.2 Hyperparameter Setting
Table 10: Comparison of success rates across different tasks in LIBERO-Spatial, evaluated on OpenVLA with INT4 quantization.
@@ 3 Ablation Study
Cost for SIS and Finetuning. SIS is a one-time process per existing expert dataset. We exploit spatio-temporal redundancy to reduce offline cost: perturbations are applied over a grid of $N\!\times\!N$ patches that evenly divide the image, and SIS is computed every $k$-th frame, reusing previous scores for intermediate frames. Empirically, $N\!=\!8$ and $k\!=\!4$ yielded robust results across all experiments and domains (Table 12). Additionally, SQIL serves as a drop-in replacement for unavoidable fine-tuning in VLA domain adaptation [2] due to variations (e.g., camera viewpoints, lighting, and hardware) using the same expert data and hyperparameters.
  k 1 2 3 4 5 6
Suc. Rate $\uparrow$ 49.3% 49.2% 49.4% 49.2% 48.2% 47.7%
Table 12: Performance comparison of timestep frequency $k$ for SIS calculation in OpenVLA with SQIL on LIBERO-Long.
               Discrepancy Metric $D$ Saliency Weight ($\beta$)       Saliency Threshold ($p$)
  Setting L2-norm KL div. 1.5 2 3 4 5 10% 15% 20% 30%
[…]
@@
Hyperparameter Robustness. We examine the sensitivity of hyperparameters within the SQIL, as detailed in Table 13. Opting for the L2-norm as the discrepancy metric $D$ yields slightly better results than the KL divergence. For SQIL, the hyperparameter $\beta$ shows stable performance within the 1.5 to 3 range. The threshold $T$ demonstrates resilience, delivering superior performance from the top 10% to 20% range, and maintaining robust results even at 30%, compared to QAT. These results confirm the robustness of our framework’s hyperparameters. As shown in Table 14, SQIL exhibits remarkable stability across hyperparameter variations, enabling shared settings across domains. Changing saliency noise distribution had negligible effect, while using raw SIS values as weights degraded performance—confirming the importance of our balanced weighting approach.